# MeshVPR: Citywide Visual Place Recognition Using 3D Meshes

### v1

@@ Abstract
Mesh-based scene representation offers a promising direction for simplifying large-scale hierarchical visual localization pipelines, combining a visual place recognition step based on global features (retrieval) and a visual localization step based on local features. While existing work demonstrates the viability of meshes for visual localization, the impact of using synthetic databases rendered from them in visual place recognition remains largely unexplored. In this work we investigate using dense 3D textured meshes for large-scale Visual Place Recognition (VPR) and identify a significant performance drop when using synthetic mesh-based databases compared to real-world images for retrieval. To address this, we propose MeshVPR, a novel VPR pipeline that utilizes a lightweight features alignment framework to bridge the gap between real-world and synthetic domains. MeshVPR leverages pre-trained VPR models and it is efficient and scalable for city-wide deployments. We introduce novel datasets with freely available 3D meshes and manually collected queries from Berlin, Paris, and Melbourne. Extensive evaluations demonstrate that MeshVPR achieves competitive performance with standard VPR pipelines, paving the way for mesh-based localization systems. Our contributions include the new task of citywide mesh-based VPR, the new benchmark datasets, MeshVPR, and a thorough analysis of open challenges. Data, code, and interactive visualizations are available at https://mesh-vpr.github.io/
@@ 1 Introduction
Estimating the location where a photo was taken based solely on its visual content is a staple of computer vision, and enables a number of useful applications ranging from augmented reality [43], robotics localization [54] and assistive technology [13]. It can be used as an alternative to GPS, where no signal nor internet connection is available or jammed [64, 12], for automatic localization of non-geotagged image and video footage, which can be useful during or after tragic events like the terror attacks of Paris in 2015, to enable a time-critical investigation instead of months of manual labour.
Different approaches have been developed to tackle this problem, depending on factors like the size of the localization map (e.g. a small building vs a large city) and the required accuracy of the estimate (e.g., a coarse position or a precise pose). In the most challenging scenarios, i.e., when the objective is the precise pose on a large map, it is common practice to rely on hierarchical solutions [18, 47, 45, 55, 56, 59, 34], which comprise two steps: (1) a Visual Place Recognition (VPR) step with global features, where methods such as NetVLAD [5] are used to obtain a coarse prediction through efficient image retrieval, and (2) a Visual Localization step with local features [46], where the initial pose estimate is refined, often by matching keypoints from the image to be localized to a 3D model of the map.
@@
figureO0.5 [Uncaptioned image] Mesh-based visual place recognition. Deploying a VPR model to a city can be performed with a database of synthetically generated images from a 3D mesh. While standard image retrieval techniques can be applied, a natural drop in results is due to the domain shift between real images (the query to be localized) and the synthetic database, requiring for the search of new solutions.
@@
Both visual place recognition and visual localization have been thoroughly studied in literature, either independently or within hierarchical pipelines. The vast majority of works relies on creating a world map with RGB images, which are used in VPR as a retrieval database, and can then be employed to create a 3D point cloud of the map, with each point being associated to a feature vector. Such point clouds are inherently tied to the model used to extract their features, limiting their flexibility and adaptability to new methods.
@@
Using a different approach, a separate branch of literature demonstrates the viability of using dense 3D textured mesh models as scene representation for visual localization [66, 51, 50, 6, 34, 35]. Overall, such methods find that the synthetic-to-real domain shift between the map and queries does not strongly impact the localization, noting that local features are robust to such visual changes. Conversely, this synthetic-to-real shift has been scarcely investigated on global features for visual place recognition: previous mesh-based localization works have either (i) used small scenes, skipping the VPR step altogether [50, 51, 6], (ii) relied on a database of real images for retrieval, before a mesh-based post-processing [34], or (iii) used retrieval on maps of limited dimension [35, 61] and did not report a noticeable performance drop related to the domain shift.
@@
Addressing this knowledge gap is crucial for enabling the creation of fully mesh-based large-scale hierarchical localization pipelines – an attractive direction due to the increasing availability of free 3D textured meshes for multiple cities [23]. Our work presents the first investigation on visual place recognition using global features on citywide 3D meshes: on a large-scale San Francisco dataset we find that a SOTA VPR model achieves a staggering 96.3% Recall@1 when using a real-images database, and drops to 76.9% when using a database made of synthetic images (see Sec. 1).
@@
To bridge this gap, we propose a novel VPR model that excels in three key aspects: (i) it efficiently matches real-world photos to a database of synthetic images, (ii) it scales effectively for large-scale datasets, enabling city-wide deployments, and (iii) it delivers strong results, achieving performance competitive with standard real-world VPR pipelines. We achieve this by developing a lightweight features alignment procedure that leverages pre-trained, state-of-the-art VPR models and adapts them to ensure consistent feature representations between real and synthetic domains. This paper details the complete pipeline, including training data preparation, features alignment, testing data acquisition from 3D meshes, and model deployment. We then evaluate our pipeline over a number of novel datasets, which rely on freely available 3D meshes and manually collected sets of queries. Finally, we demonstrate how our pipeline, called MeshVPR, allows to achieve impressive results on each of the datasets, performing competitively with standard VPR pipelines using real-world database images. To summarize, our contributions are:
@@
  • A thorough analysis of failure cases and open challenges, noting that despite the large domain gap MeshVPR is able to achieve astonishing quantitative and qualitative results, leading the way for future research and for new mesh-based localization systems.
We encourage readers to check our webpage [^1] https://mesh-vpr.github.io/ with explanatory videos, visualizations, code, datasets and trained models.
@@ 2 Related work
Synthetic data for localization tasks Despite the existence of numerous synthetic 3D meshes of cities [14, 23] and synthetically-generated environments [1, 25, 24, 40, 39, 49, 15], their usage for image localization has been little explored in recent years. Among the few outliers, Panek et al [34, 35] proposes the use of 3D meshes for visual localization, although still relying on real-world images for the retrieval step. On a separate development Vallone et al [61] spearheaded the exploration of aerial images to render street-view images for place recognition, and used it to generate a database of 44k images for the coarse localization of real-world query images.
Taking such literature into account, this paper fills a noticeable gap, as we are the first to (i) explore 3D meshes for citywide VPR, (ii) quantify the gap in results between using real images and synthetic images, (iii) use features alignment to efficiently adopt robust SOTA VPR models and (iv) provide a number of large-scale datasets comprising challenging manually-collected query images.
@@ 3 Mesh-based Visual Place Recognition Pipeline
Motivated by the ever-growing availability of freely-available citywide 3D textured mesh models (hereinafter simply referred to as 3D meshes) [23], and the recent advances in mesh-based visual localization [34, 35], our idea is to use such 3D meshes to obtain a synthetic database.
Formally, given a target city where we want to deploy a VPR system, we define 3D mesh as $\mathcal{M}_{target}$, and the database of images extracted from it as $(\mathcal{X}_{target}^{synt},\mathcal{Y}_{target})$, the goal is to correctly match a given query image to a synthetic image from $\mathcal{X}_{target}^{synt}$ that depicts the same scene. To allow VPR models to perform well in such conditions (i.e. real query against a synthetic database), we devised a new mesh-based VPR pipeline that can be applied to any existing VPR model $f_{\theta_{real}}$, trained on real-world data. Due to the appearance gap between real and synthetic images (see Fig. 1), the features that $f_{\theta_{real}}$ extracts from a couple of real/synthetic images, even from the same viewpoint, are not well aligned. To address this problem, we use a new model $f_{\theta_{synt}}$, that is initialized by setting $\theta_{synt}=\theta_{real}$ and then fine-tuned so that the embeddings $f_{\theta_{synt}}(x_{synt})$ and $f_{\theta_{real}}(x_{real})$ for a pair of matching synthetic/real images (i.e., from the same viewpoint) are well aligned. This is akin to a teacher-student paradigm, where $f_{\theta_{real}}$ is the teacher and $f_{\theta_{synt}}$, although this differs with many teacher-student models in that (i) the two models have the same architecture and (ii) they are trained on images from different domains that share the exact same viewpoint. Finally, the deployment is completed by using the specialized model $f_{\theta_{synt}}$ to extract the embeddings from the synthetic database $\mathcal{X}_{target}^{synt}$, whereas the model $f_{\theta_{real}}$ extracts the embeddings from the real-world queries. This idea is illustrated in Sec. 3 (right).
@@
The overall deployment pipeline is a 5-steps recipe, as illustrated in Fig. 2. Remarkably, the features alignment (steps 1-3) needs to be performed only once to align the model $f_{\theta_{synt}}$, while any actual deployment only requires the database extraction from the 3D mesh of the target city.
@@ 3.2 Step 2: Generate alignment images from mesh
After having downloaded the citywide 3D mesh and the corresponding training dataset, it is then possible to create a set of synthetic images $\mathcal{X}_{align}^{synt}$ from the mesh $\mathcal{M}_{align}$, such that they precisely mirror the images from $\mathcal{X}_{align}^{real}$ (see Fig. 1). This requires to have the latitude, longitude and heading (yaw) of each real image in $\mathcal{X}_{align}^{real}$, as they are needed to create views of the mesh from the corresponding viewpoints. Altitude, pitch and roll are also required to accurately replicate the images, but these can be inferred from the physics information of the 3D mesh. To accomplish that, we cast a line trace from the real image location towards the ground and see where it intersects with the model. This intersection point serves as a reliable estimate of the ground level. The views are then generated at a height of 2.5m, i.e., roughly the height of a typical car-mounted camera used to collect VPR datasets [31, 28]. Additionally, the normal vector at the intersection is used to estimate pitch and roll.
Given a 6 DoF camera pose, its synthetic counterpart can be generated from the mesh using any real-time rendering engine: in our case, we rely on Unreal Engine and Cesium.
@@ 3.3 Step 3: Features alignment
To this end, we initialize $f_{\theta_{real}}=f_{\theta_{synt}}$ to any open-source pretrained VPR model, like CosPlace [8], MixVPR [3] or SALAD [19]. Secondly, we fine-tune $f_{\theta_{synt}}$ on the images $\mathcal{X}_{align}^{synt}$ to mirror the features extracted by $f_{\theta_{real}}$ from the corresponding images in $\mathcal{X}_{align}^{real}$. More formally, we optimize the following MSE loss:
@@
Let us further highlight that: (i) the training phase (steps 1-3) needs to be performed only once to create the model $f(\theta_{synt})$, which can then be deployed to any new city, and (ii) although these steps do require a real-world dataset $\mathcal{X}_{align}^{real}$ annotated with position and heading, it is possible to use existing datasets such as SF-XL [8] without having to rely on images from the target city.
@@ 3.4 Step 4: Generate the test database
The next step consists in generating the database $(\mathcal{X}_{target}^{synt},\mathcal{Y}_{target})$ for the target city(ies), i.e., the one(s) where we want to deploy the VPR system. Given the 3D mesh $\mathcal{M}_{target}$ from the target city, we want to ensure that the database is constructed by generating relevant views, which could ideally match the camera viewpoints of the queries.
We generate street-view-like images by simulating in the 3D mesh the path that a camera-equipped car would take while driving on the streets to collect a database. This process begins with fetching OpenStreetMap (OSM) street data for the area covered by the 3D mesh $\mathcal{M}_{target}$. Then, a graph is constructed from the OSM data, and a path that visits every edge is computed by solving the Route Inspection Problem. This continuous path is needed in order to ensure equal spacing between the individual sampling locations. The actual capture is performed by moving the camera along the path by a configurable distance (10 meters in our experiments, unless otherwise stated).
@@ 4 Test sets
To empirically validate the soundness of MeshVPR, we built three test sets, each one consisting of a synthetic geo-tagged database, built following the method in Sec. 3.4, and a number of real-world queries. The three test sets cover the cities of Berlin, Paris and Melbourne, for which 3D meshes are freely available: the datasets characteristics are summarized in Tab. 1. For Berlin and Paris, a part of the queries have been manually collected, and another part has been downloaded from Flickr and Wikimedia, which are commonly used sources for computer vision datasets [37, 38, 8]. Some examples are shown in the qualitative results in Sec. 5 and in the Supplementary. Given the inaccuracies in GPS positioning (both in our manually collected queries, and Flickr and Wikimedia photos), we manually selected only a few hundred queries for the two European capitals, carefully removing any image for which the GPS label did not match its actual position. Using a different approach, for Melbourne we use as queries a subset of the Mapillary Street-Level Sequences dataset [63] collected in the city of Melbourne.
@@ 5.1 Implementation details
Features Alignment. Our feature alignment is performed by pairing two copies of the same VPR model (e.g., NetVLAD, CosPlace, MixVPR), both initialized with open-source pretrained weights. One of the two models, which we refer to as $f_{\theta_{real}}$, has frozen weights $\theta_{real}$, while the weights $\theta_{synt}$ of $f_{\theta_{synt}}$ are fine-tuned for 50k iterations with a batch size of 32 and the Adam [22] optimizer with learning rate $51e-$ 5. As per the training set, we use the largest publicly available VPR dataset, namely San Francisco eXtra Large (SF-XL) [8] and its 3D mesh counterpart, with synthetic views extracted from the San Francisco HQ mesh (see Tab. 1).
Inference. At inference time, $f_{\theta_{synt}}$ is used to extract features from synthetic database, while the queries are processed through $f_{\theta_{real}}$. As in standard VPR [5, 16, 2, 8, 3, 10], queries’ features are matched against database ones by kNN algorithm (see Sec. 3, right).

### v2

@@ Abstract
Mesh-based scene representation offers a promising direction for simplifying large-scale hierarchical visual localization pipelines, combining a visual place recognition step based on global features (retrieval) and a visual localization step based on local features. While existing work demonstrates the viability of meshes for visual localization, the impact of using synthetic databases rendered from them in visual place recognition remains largely unexplored. In this work we investigate using dense 3D textured meshes for large-scale Visual Place Recognition (VPR). We identify a significant performance drop when using synthetic mesh-based image databases compared to real-world images for retrieval. To address this, we propose MeshVPR, a novel VPR pipeline that utilizes a lightweight features alignment framework to bridge the gap between real-world and synthetic domains. MeshVPR leverages pre-trained VPR models and is efficient and scalable for city-wide deployments. We introduce novel datasets with freely available 3D meshes and manually collected queries from Berlin, Paris, and Melbourne. Extensive evaluations demonstrate that MeshVPR achieves competitive performance with standard VPR pipelines, paving the way for mesh-based localization systems. Data, code, and interactive visualizations are available at https://meshvpr.github.io/
@@ 1 Introduction
Estimating the location of where a photo was taken based solely on its visual content is a staple of computer vision, and enables a number of applications ranging from augmented reality [45], robotics localization [56] and assistive technology [13]. It can be used as an alternative to GPS, where no signal nor internet connection is available or jammed [68, 12]. Additionally, it can help with automatic localization of non-geotagged image and video footage, which can be useful during or after tragic events like the terror attacks of Paris in 2015, to enable a time-critical investigation instead of months of manual labour.
The approaches developed to tackle the problem depend on factors like the size of the localization map (e.g. a small building vs a large city) and the required accuracy of the estimate (e.g., a coarse position or a precise pose). In the most challenging scenarios, i.e., when the objective is the precise pose on a large map, it is common practice to rely on hierarchical solutions [18, 49, 47, 57, 59, 63, 36], which comprise two steps: (1) a Visual Place Recognition (VPR) step with global features, where methods such as NetVLAD [5] are used to obtain a coarse prediction through efficient image retrieval, and (2) a Visual Localization (VL) step with local features [48], where the initial pose estimate is refined, often by matching keypoints from the image to be localized to a 3D model of the map.
@@
figureO0.5 [Uncaptioned image] Mesh-based VPR. Deploying a VPR model to a city can be performed with a database of synthetically generated images from a 3D mesh. While standard image retrieval techniques can be applied, a natural drop in results is due to the domain shift between real images (the query to be localized) and the synthetic database, requiring for the search of new solutions.
@@
Both VPR and VL have been thoroughly studied in literature, either independently or within hierarchical pipelines. The vast majority of works relies on creating a world map with RGB images, which are used in VPR as a retrieval database, and can then be employed to create a 3D point cloud of the map, with each point being associated to a feature vector. Such point clouds are inherently tied to the model used to extract their features, limiting their flexibility and adaptability to new methods.
@@
Using a different approach, a separate branch of literature demonstrates the viability of using dense 3D textured mesh models as scene representation for visual localization [70, 53, 52, 6, 36, 37]. Overall, such methods find that the synthetic-to-real domain shift between the map and queries does not strongly impact the localization, noting that local features are robust to such visual changes. Conversely, this synthetic-to-real shift has been scarcely investigated on global features for VPR: previous mesh-based localization works have either (i) used small scenes, skipping the VPR step altogether [52, 53, 6], (ii) relied on a database of real images for retrieval, before a mesh-based post-processing [36], or (iii) used retrieval on maps of limited dimension [37, 65] and did not report a noticeable performance drop related to the domain shift.
@@
Addressing this knowledge gap is crucial for enabling the creation of fully mesh-based large-scale hierarchical localization pipelines – an attractive direction due to the increasing availability of free 3D textured meshes for multiple cities [23]. This work is the first investigation on VPR using global features on citywide 3D meshes: on a large-scale San Francisco dataset we find that a SOTA VPR model achieves a Recall@1 of 96.3% when using a real-images database, and drops to 76.9% when using a database made of synthetic images (see Sec. 1).
@@
To bridge this gap, we propose a novel VPR model that excels in three key aspects: (i) it efficiently matches real-world photos to a database of synthetic images, (ii) it scales effectively for large-scale datasets, enabling city-wide deployments, and (iii) it delivers strong results, achieving performance competitive with standard real-world VPR pipelines. We achieve this by developing a lightweight features alignment procedure that leverages pre-trained, state-of-the-art VPR models and adapts them to ensure consistent feature representations between real and synthetic domains. This paper details the complete pipeline, including training data preparation, features alignment, testing data acquisition from 3D meshes, and model deployment. We then evaluate our pipeline over a number of novel datasets, which rely on freely available 3D meshes and manually collected sets of queries. Finally, we demonstrate how our pipeline, called MeshVPR, achieves impressive results on each of the datasets, performing competitively with standard VPR pipelines using real-world database images. To summarize, our contributions are:
@@
  • A thorough analysis of failure cases and open challenges, noting that despite the large domain gap MeshVPR is able to achieve excellent quantitative and qualitative results, leading the way for future research and for new mesh-based localization systems.
@@ 2 Related work
Synthetic data for localization tasks The use of synthetic images has been explored for pose estimation with respect to a large indoor 3D map [58]. Aiming at addressing changes in appearances, [61] has explored view synthesis to align viewpoints of queries and database. Despite the existence of numerous synthetic 3D meshes of cities [14, 23] and synthetically-generated environments [1, 25, 24, 42, 41, 51, 15], their usage for image localization has been little explored in recent years. Among the few outliers, Panek et al [36, 37] proposes the use of 3D meshes for visual localization, although still relying on real-world images for the retrieval step. On a separate development, Vallone et al [65] spearheaded the exploration of aerial images to render street-view images for place recognition, and used it to generate a database of 44k images for the coarse localization of real-world query images. More in general, [33] explored the use of synthetic data to improve the localization process, while others further train the model with real and synthetic data to improve feature representation [32].
This paper fills a noticeable gap in the literature, as it is the first work to (i) explore 3D meshes for citywide VPR, (ii) quantify the gap in results between using real images and synthetic images, (iii) use features alignment to efficiently adopt robust SOTA VPR models and (iv) provide a number of large-scale datasets comprising challenging manually-collected query images.
@@ 3 Mesh-based Visual Place Recognition Pipeline
Motivated by the ever-growing availability of freely-available citywide 3D textured mesh models [23], and the recent advances in mesh-based visual localization [36, 37], our idea is to use such 3D meshes to obtain a synthetic database. Note that these meshes are generally built from aerial images, which can not directly be used for VPR. Formally, given a target city where we want to deploy a VPR system, we define 3D mesh as $\mathcal{M}_{target}$, and the database of images extracted from it as $(\mathcal{X}_{target}^{synt},\mathcal{Y}_{target})$, the goal is to correctly match a given query image to a synthetic image from $\mathcal{X}_{target}^{synt}$ that depicts the same scene. To allow VPR models to perform well in such conditions (i.e. real query against a synthetic database), we devised a new mesh-based VPR pipeline that can be applied to any existing VPR model $f_{\theta_{real}}$, trained on real-world data. Due to the appearance gap between real and synthetic images (see Fig. 1), the features that $f_{\theta_{real}}$ extracts from a couple of real/synthetic images, even from the same viewpoint, are not well aligned. To address this problem, we use a new model $f_{\theta_{synt}}$, that is initialized by setting $\theta_{synt}=\theta_{real}$ and then fine-tuned so that the embeddings $f_{\theta_{synt}}(x_{synt})$ and $f_{\theta_{real}}(x_{real})$ for a pair of matching synthetic/real images (i.e., from the same viewpoint) are well aligned. This is akin to a teacher-student paradigm, where $f_{\theta_{real}}$ is the teacher and $f_{\theta_{synt}}$, although this differs with many teacher-student models in that (i) the two models have the same architecture and (ii) they are trained on images from different domains that share the exact same viewpoint. Finally, the deployment is completed by using the specialized model $f_{\theta_{synt}}$ to extract the embeddings from the synthetic database $\mathcal{X}_{target}^{synt}$, whereas the model $f_{\theta_{real}}$ extracts the embeddings from the real-world queries. This idea is illustrated in Sec. 3 (right). The overall deployment pipeline is a 5-steps recipe, as illustrated in Fig. 2. Remarkably, the features alignment (steps 1-3) needs to be performed only once to align the model $f_{\theta_{synt}}$, while any actual deployment only requires the database extraction from the 3D mesh of the target city.
@@ 3.2 Step 2: Generate alignment images from mesh
With downloaded the citywide 3D mesh and the corresponding training dataset, it is then possible to create a set of synthetic images $\mathcal{X}_{align}^{synt}$ from the mesh $\mathcal{M}_{align}$, such that they precisely mirror the images from $\mathcal{X}_{align}^{real}$ (see Fig. 1).
This requires to have the latitude, longitude and heading (yaw) of each real image in $\mathcal{X}_{align}^{real}$, as they are needed to create views of the mesh from the corresponding viewpoints. Altitude, pitch and roll are also required to accurately replicate the images, but these can be inferred from the physics information of the 3D mesh. To accomplish that, we cast a ray from the real image location towards the ground and see where it intersects with the model. This intersection point serves as a reliable estimate of the ground level. The views are then generated at a height of 2.5m, i.e., roughly the height of a typical car-mounted camera used to collect VPR datasets [31, 28]. Additionally, the normal vector at the intersection is used to estimate pitch and roll.
Synthetic counterpart for real images can be generated from the mesh using any rendering engine: in our case, we rely on Unreal Engine and Cesium.
@@ 3.3 Step 3: Features alignment
To this end, we initialize $f_{\theta_{real}}=f_{\theta_{synt}}$ to any open-source pretrained VPR model, like CosPlace [8], MixVPR [3] or SALAD [19]. Secondly, we fine-tune $f_{\theta_{synt}}$ on the images $\mathcal{X}_{align}^{synt}$ to mirror the features extracted by $f_{\theta_{real}}$ from the corresponding images in $\mathcal{X}_{align}^{real}$. Formally, we optimize the following MSE loss:
@@
Note that: (i) the training phase (steps 1-3) needs to be performed only once to create the model $f(\theta_{synt})$, which can then be deployed to any new city, and (ii) although these steps do require a real-world dataset $\mathcal{X}_{align}^{real}$ annotated with position and heading, it is possible to use existing datasets such as SF-XL [8] without having to rely on images from the target city.
@@ 3.4 Step 4: Generate the test database
The next step consists in generating the database $(\mathcal{X}_{target}^{synt},\mathcal{Y}_{target})$ for the target cities, i.e., the ones where we want to deploy the VPR system. Given the 3D mesh $\mathcal{M}_{target}$ from the target city, we want to ensure that the database is constructed by generating relevant views, which could ideally match the camera viewpoints of the queries.
We generate street-view-like images from the 3D mesh by simulating the path that a camera-equipped car would take to collect a database. This process begins with fetching OpenStreetMap (OSM) street data for the area covered by the 3D mesh $\mathcal{M}_{target}$. Then, a graph is constructed from the OSM data. For this graph we compute a path that visits every edge, this is accomplished by solving the Route Inspection Problem for the graph. This continuous path is needed in order to ensure equal spacing between the individual sampling locations and reduce the sampling time. The actual capture is performed by moving the camera along the path by a configurable distance (10 meters in our experiments, unless otherwise stated). We note that sampling can easily be adapted to the specific application, e.g. a database for drone localization can be sampled from simulated aerial viewpoints. This underscores the flexibility of our approach for scenarios that are different from street-level visual place recognition.
@@ 4 Test sets
To empirically validate the soundness of MeshVPR, we built three test sets, each one consisting of a synthetic geo-tagged database, built following the method in Sec. 3.4, and a number of real-world queries. The three test sets cover the cities of Berlin, Paris and Melbourne, for which 3D meshes are freely available: the datasets characteristics are summarized in Tab. 1. For Berlin and Paris, a part of the queries have been manually collected, and another part has been downloaded from Flickr and Wikimedia, which are commonly used sources for computer vision datasets [39, 40, 8]. Some examples are shown in the qualitative results in Sec. 5 and in the Supplementary. Given the inaccuracies in GPS positioning (both in our manually collected queries, and Flickr and Wikimedia photos), we manually selected only a few hundred queries for the two European capitals, carefully removing any image for which the GPS label did not match its actual position. For Melbourne we use as queries a subset of the Mapillary Street-Level Sequences dataset [67] collected in the city of Melbourne.
@@ 5.1 Implementation details
Features Alignment. Our feature alignment is performed by pairing two copies of the same VPR model (e.g., NetVLAD, CosPlace, MixVPR), both initialized with open-source pretrained weights. One of the two models, which we refer to as $f_{\theta_{real}}$, has frozen weights $\theta_{real}$, while the weights $\theta_{synt}$ of $f_{\theta_{synt}}$ are fine-tuned for 50k iterations with a batch size of 32 and the Adam [22] optimizer with learning rate $51e-$ 5. Training takes only 3 hours on just 4GB of VRAM when using the heaviest model considered (i.e. SALAD) on a Nvidia RTX-4090 GPU.
As per the training set, we use the largest publicly available VPR dataset, namely San Francisco eXtra Large (SF-XL) [8] and its 3D mesh counterpart, with synthetic views extracted from the San Francisco HQ mesh (see Tab. 1) unless otherwise stated.
Inference. At inference time, $f_{\theta_{synt}}$ is used to extract features from synthetic database, while the queries are processed through $f_{\theta_{real}}$. As in standard VPR [5, 16, 2, 8, 3, 10], queries’ features are matched against database ones by kNN algorithm (see Sec. 3, right). We use a threshold of 100 meters for positives.